OpenStudy (anonymous):

the volume of a box F is given by the formula v=lwh, where l is the lenght, w is the width, and h is the height. a. Solve the formula for h. b. what is the height of a box with a volume of 50 cubic meters , lenght of 10 meters, and widht of 2 meters?

OpenStudy (prizzyjade):

h=v/lw h=50cubic meters/(10m*2m) h=2.5meters

OpenStudy (anonymous):

do you have to multiply all of them

OpenStudy (prizzyjade):

Yeah since volume = l*w*h